OpenStudy (anonymous):

Find the point-slope form of the line that satisfies the following: Slope 4, passing through (-3,1).

OpenStudy (anonymous):

do you know slope intercept form?

OpenStudy (anonymous):

what formulas do you know for linear algebra?

OpenStudy (anonymous):

What is the point-slope form of a line? Find the point-slope form of the line that satisfies the following: Slope 4, passing through (-3,1). that how it looks

OpenStudy (anonymous):

this is point slope form y - y1 = m ( x - x1 ) slope y - y1 m = -------- x - x1

OpenStudy (anonymous):

y - y1 = m ( x - x1 ) ^ slope

OpenStudy (theeric):

Adding to mth3v4's response, x1 and y1 are the x and y values of a point on the line. The point you have is (-3, 1). So, to have the line written in that form, you would just substitute in your values.

OpenStudy (anonymous):

Provide the class with two problems. The first should provide a point and a slope. The second should provide two points. For each, have your classmates find: 1) the point-slope form of the line and 2) the slope-intercept form of the line. Also, state the value of the slope and intercept for each.

OpenStudy (theeric):

That should probably go up as another question! Haha!

OpenStudy (anonymous):

I hate math that last thing I sent you was the last bit of my discussion question